OpenStudy (anonymous):

Suppose that f is a function with f(-2)=-2 and f'(-2)=32. If h(x)=(1/(f(x)^5)) then h'(-2)=?

OpenStudy (anonymous):

Use the chain rule: \[h(x)=f(x)^{-5}\] \[h'(x) = -5f(x)^{-6}f'(x)\]

OpenStudy (anonymous):

Now just plug in -2: h'(-2) = -5/(f(-2)^6) * f'(-2) = -5/64 * 32 = -5/2